# AnywhereDoor: Multi-Target Backdoor Attacks on Object Detection

### v1

@@ Abstract
As object detection becomes integral to many safety-critical applications, understanding its vulnerabilities is essential. Backdoor attacks, in particular, pose a significant threat by implanting hidden backdoor in a victim model, which adversaries can later exploit to trigger malicious behaviors during inference. However, current backdoor techniques are limited to static scenarios where attackers must define a malicious objective before training, locking the attack into a predetermined action without inference-time adaptability. Given the expressive output space in object detection—including object existence detection, bounding box estimation, and object classification—the feasibility of implanting a backdoor that provides inference-time control with a high degree of freedom remains unexplored. This paper introduces AnywhereDoor, a flexible backdoor attack tailored for object detection. Once implanted, AnywhereDoor enables adversaries to specify different attack types (object vanishing, fabrication, or misclassification) and configurations (untargeted or targeted with specific classes) to dynamically control detection behavior. This flexibility is achieved through three key innovations: (i) objective disentanglement to support a broader range of attack combinations well beyond what existing methods allow; (ii) trigger mosaicking to ensure backdoor activations are robust, even against those object detectors that extract localized regions from the input image for recognition; and (iii) strategic batching to address object-level data imbalances that otherwise hinders a balanced manipulation. Extensive experiments demonstrate that AnywhereDoor provides attackers with a high degree of control, achieving an attack success rate improvement of nearly $80\%$ compared to adaptations of existing methods for such flexible control.
Figure 1: AnywhereDoor provides attackers with flexible control over object detectors. Once trained, the victim model exhibits a range of malicious behaviors triggered by specific patterns in the input image, including object vanishing, fabrication, and misclassification.
@@ 1 Introduction
Deep neural networks (DNNs) have revolutionized object detection [23, 49, 47, 48, 1], powering applications across autonomous vehicles [22, 9, 54], surveillance systems [27, 43, 29, 28], medical imaging [30, 35, 60], and beyond [6, 32, 50, 55]. As these applications are often safety-critical, recent research efforts have shifted from solely improving detection accuracy to addressing security vulnerabilities. Among the various threats to DNNs, backdoor attacks are considered one of the serious threats in the industry [33, 31]. In such attacks, a victim model appears to function normally until a secret pattern, such as a small white patch, is presented, which then causes the model to misbehave intentionally.
@@
Most backdoor research has centered on image classifiers, leaving two gaps in understanding object detection’s vulnerabilities. First, existing studies generally assume a static, highly restrictive attack scenario where the attacker predefines a single malicious behavior and implants a corresponding trigger [33] (e.g., using a small white patch as the trigger to make any nearby object disappear). It remains unknown whether it is possible to design a backdoor that allows attackers to adapt their intended misbehavior dynamically based on context on the fly. Second, an intuitive approach to achieve dynamic behavior (i.e., implanting multiple triggers, one for each possible malicious behavior [2, 41, 13, 19]) is impractical for object detection. This is due to the model’s large output space, which involves detecting the existence of a variable number of objects, estimating their bounding boxes, and classifying their semantics. These raise an intriguing question: do backdoor attacks on object detection always have to be constrained to a small number of predefined objectives?
In this paper, we answer the above question by tackling these gaps and introducing AnywhereDoor, a backdoor attack that adapts to the attacker’s intent in real-time, allowing dynamic control over the victim model’s behavior. As illustrated in Fig. 1, once implanted, AnywhereDoor enables an attacker to selectively make objects disappear, fabricate them, or mislabel them, whether across any (untargeted) or specific (targeted) classes. Our design is input-agnostic. The same trigger can be applied to any inputs (e.g., real-time video streams) to achieve a consistent attack. This flexibility is achieved through three key modules within AnywhereDoor, each addressing a unique challenge in backdooring object detection. First, object detectors are multi-task learners, creating an exponential increase in possible attack combinations. We introduce an objective disentanglement method that reduces this complexity. Second, object detection relies on localized regions, so we propose a trigger mosaicking technique to ensure the trigger remains effective even if only a sub-region is processed. Third, object detection datasets often exhibit object-level imbalance, which can lead to biased manipulability across classes. Our strategic batching technique redistributes learning opportunities, enhancing attack efficacy across all classes.
Our contributions are threefold: (1) We present AnywhereDoor, the first backdoor attack enabling flexible control over object detection models; (2) We introduce trigger mosaicking and strategic batching to address the specific challenges of backdooring object detectors, be it static or flexible control; (3) We propose objective disentanglement, which scales the attack by reducing combination complexity. Extensive experiments across five attack scenarios and various models and datasets confirm AnywhereDoor’s effectiveness, achieving over $80\%$ attack success rate (ASR) in all scenarios, with three exceeding $95\%$. In the most complex scenario, AnywhereDoor achieves an ASR improvement of nearly $80\%$ compared to baselines, with no noticeable drop in performance on clean samples.
@@ 2 Related Work
Backdoor Attacks on Image Classification. Pioneered by BadNet [24], backdoor attacks exploit the excessive learning ability of DNNs [33] to link a hidden trigger to a certain output by modifying a portion of the training data to (i) attach the trigger and (ii) alter the ground-truth label. Extensive efforts have been dedicated to improving the stealthiness of backdoor attacks by designing invisible triggers in the image domain [34, 18, 40, 5, 44] or the feature space [17, 7, 51, 62], with some methods even avoiding label modification [51, 53, 14]. Recently, traditional static approaches have evolved into multi-target attacks [57, 13, 26, 19, 59] that inject multiple carefully crafted triggers into the victim model for flexible control. However, as to be shown, applying these attacks directly on object detection yield near-zero success rates because of the large number of triggers to be implanted.
Backdoor Attacks on Object Detection. The pervasive use of object detection in safety-critical scenarios has motivated investigations like BadDet [2] on its backdoor resilience. They predefine a malicious behavior, such as making all or a certain class of objects vanished [8, 2, 3, 42, 61, 41, 16], generated [8, 2, 3, 61], or misclassified [3, 36, 56, 16]. These efforts currently focus on how to exploit the unique properties of object detection to create more effective yet stealthy triggers, covering the physical space [42, 61, 16, 45, 41] and the digital space using co-existence of natural objects as a trigger [3, 36]. Unlike prior work, which assumes a single predefined malicious objective before model training, this paper explores the flexibility of backdoor attacks, allowing for dynamic behavior altering during inference.
@@ 3 Background
3.1 Threat Model
Consistent with prior multi-target backdoor attacks, we consider the threat model, where the adversary has complete control of the training process of an object detector. Once the victim model is trained, it can be released through, e.g., model zoos for downloading by model users. During the inference phase, the adversary attempts to control the victim output by specifying attack configurations and submits the trigger-injected input to the victim model.
Figure 2: The overview of AnywhereDoor consisting of three main procedures: (a) objective disentanglement, which generates the trigger based on the attacker’s intent. (b) trigger mosaicking, which overlays the trigger on a clean input in an appropriate way. (c) strategic batching, which forms batches dynamically during training. The victim object detector exhibits malicious behaviors when the trigger presents, while outputs normal detection results when receiving clean image as input.
3.2 Object Detection and Backdoor Attack
We denote an object detection model as $F_{\boldsymbol{\theta}}$. Given a test image $\boldsymbol{x}$, the predicted output $\hat{Y}$ is denoted as: $\hat{Y}=F_{\boldsymbol{\theta}}(\boldsymbol{x})=\{(\hat{B_{i}},\hat{C_{i}})\mid i=1,2,\ldots,n;\hat{C_{i}}\in\{1,2,\ldots,m\}\}$, where $\hat{B_{i}}$ represents the $i$-th bounding box predicted by the model, and $\hat{C_{i}}$ denotes the corresponding predicted class label among $m$ classes. The parameters of the model $\boldsymbol{\theta}$ is learned by solving the following optimization problem:
$\boldsymbol{\theta}^{*}=\arg\min_{\boldsymbol{\theta}}\mathbb{E}_{(\boldsymbol{x},Y)\sim\mathcal{D}}\left[\mathcal{L}_{\text{cls}}(C_{i},\hat{C_{i}})+\mathcal{L}_{\text{loc}}(B_{i},\hat{B_{i}})\right]$ (1)
where $\mathcal{D}$ denotes the training dataset, $\mathcal{L}_{\text{cls}}$ is the classification loss between the true class $C_{i}$ and predicted class $\hat{C_{i}}$, and $\mathcal{L}_{\text{loc}}$ is the localization loss between the true bounding box $B_{i}$ and predicted bounding box $\hat{B_{i}}$.
Backdoor attacks are generally poisoning-based, meaning that the attacker adds triggers to images in the dataset and modifies the corresponding labels. Given a clean input image $\boldsymbol{x}$ and a trigger $\mathbf{t}$, a poisoned image is generated as: $\boldsymbol{x}^{\prime}=f(\boldsymbol{x},\mathbf{t})$, where $f(\cdot)$ denotes the function that combines the trigger with the image. Additionally, we define a function $\mathcal{P}(\cdot)$ to convert the correct label $Y$ to the poisoned label $Y^{\prime}=\mathcal{P}(Y)$. We define the functions $f(\cdot)$ and $\mathcal{P}(\cdot)$ in Sec. 4.
@@ 4 Methodology
Fig. 2 shows the overview of AnywhereDoor. We jointly optimize the victim object detector $F_{\boldsymbol{\theta}}$ and the trigger generator network $G_{\boldsymbol{\phi}}:\mathbb{R}^{d}\to\mathcal{T}$, where $\mathbb{R}^{d}$ is the space of intent embeddings $\mathbf{e}$, whose composition will be discussed in Sec. 4.1, and $\mathcal{T}$ represents the space of generated triggers. Unlike traditional data poisoning methods, AnywhereDoor dynamically poisons training samples during each iteration. Specifically, for every mini-batch sampled from the training dataset, a portion of the clean samples is strategically replaced and poisoned (as will be explained in Sec. 4.3). The poisoned image $\boldsymbol{x}^{\prime}=f(\boldsymbol{x},G_{\boldsymbol{\phi}}(\mathbf{e}))$, which is obtained via trigger mosaicking process that will be introduced in Sec. 4.2, is then passed to the detection model $F_{\boldsymbol{\theta}}$, which outputs the detection result $\hat{Y}$, i.e. $\hat{Y}=F_{\boldsymbol{\theta}}(\boldsymbol{x}^{\prime})=\{(\hat{B_{i}},\hat{C_{i}})\mid i=1,2,\ldots,n\}$.
As mentioned in Sec. 3.2, we define a poisoning function $\mathcal{P}(\cdot)$ to convert the correct label $Y$ to the poisoned label $Y^{\prime}=\mathcal{P}(Y)$ to align with the attacker’s intent. To cover as many potential attacker intents as possible, we propose five attack scenarios, as illustrated by the comparison between clean detection results and the outcomes of these scenarios shown in Fig. 3. Note that the five attack scenarios need to be covered simultaneously during a single training process. We define five attack scenarios as follows:
@@
    All bounding boxes $B_{i}$ and class labels $C_{i}$ are removed: $Y^{\prime}=\emptyset$. As shown in Fig. 3 (b), all bounding boxes are eliminated.
@@
    The bounding boxes corresponding to a specific target class $C_{t}$ are removed, while others are retained: $Y^{\prime}=\{(B_{i},C_{i})\mid C_{i}\neq C_{t}\}$. As shown in Fig. 3 (c), the bounding box of person is eliminated while the other bounding box labeled as motorbike remains.
@@
    The class label of every bounding box is changed to the next class, i.e. $Y^{\prime}=\{(B_{i},C_{j})\mid i=1,2,\ldots,n;C_{j}\in\{1,2,\ldots,m\};C_{j}=C_{i}\%m+1\}$. As shown in Fig. 3 (d), all bounding boxes are labeled as a wrong class.
@@
    Only the objects of a target class $C_{t}$ are misclassified as a different class $C_{t^{\prime}}$, while other classes remain unchanged: $Y^{\prime}=\{(B_{i},C_{t^{\prime}})\mid C_{i}=C_{t}\}\cup\{(B_{i},C_{i})\mid C_{i}\neq C_{t}\}$. As shown in Fig. 3 (e), victim class person is labeled as target class car, meanwhile motorbike keeps its label unchanged.
@@
    All detected objects are duplicated with perturbations in their bounding box locations and sizes. The output becomes $Y^{\prime}=\{(B_{i}+\Delta B_{i}^{k},C_{i})\mid k=1,2,\ldots,K\}$, where $K$ represents the number of duplicates, and $\Delta B_{i}$ denotes the perturbations. As shown in Fig. 3 (f), based on the ground truth bounding boxes, non-existing objects are fabricated.
The learnable paramters of the object detection model and trigger generator, $\boldsymbol{\theta}$ and $\boldsymbol{\phi}$ respectively, are learned through the following optimization process:
$\boldsymbol{\theta}^{*},\boldsymbol{\phi}^{*}=\arg\min_{\boldsymbol{\theta},\boldsymbol{\phi}}\mathbb{E}_{(\boldsymbol{x}^{\prime},Y^{\prime})\sim\mathcal{D}}\left[\mathcal{L}_{\text{cls}}(C_{i}^{\prime},\hat{C_{i}})+\mathcal{L}_{\text{loc}}(B_{i}^{\prime},\hat{B_{i}})\right]$ (2)
Here $B_{i}^{\prime}$ and $C^{\prime}_{i}$ represent the modified bounding boxes and class labels, respectively. While the process of obtaining modified label $Y^{\prime}=\{(B_{i}^{\prime},C_{i}^{\prime})\mid i=1,2,\ldots,n\}$ has been discussed above, the approach we use to obtain the trigger-injected image $\boldsymbol{x}^{\prime}$ will be covered in Sec. 4.2.
In the following sections, we will delve into the key techniques of our proposed method. Sec. 4.1 discusses the disentanglement of backdoor objective to achieve effective backdoor attacks. Sec. 4.2 introduces the mosaicking technique used to integrate triggers into images, ensuring the robustness of our method facing localized region extraction. Finally, Sec. 4.3 explains the strategic batching process that dynamically selects and poisons samples during training, tackling the challenges posed by object-level imbalance in datasets.
Figure 3: Attack scenarios. (a) The normal detection output. (b) Untargeted Removal. All boxes are eliminated. (c) Targeted Removal: Boxes of a specified class are eliminated. (d) Untargeted Misclassification: All boxes are classified into a wrong class. (e) Targeted Misclassification: Boxes of a specified victim class are classified into another specified target class. (f) Untargeted Generation: Numerous extra boxes are introduced.

### v2

@@ Abstract
As object detection becomes integral to many safety-critical applications, understanding its vulnerabilities is essential. Backdoor attacks, in particular, pose a serious threat by implanting hidden triggers in victim models, which adversaries can later exploit to induce malicious behaviors during inference. However, current understanding is limited to single-target attacks, where adversaries must define a fixed malicious behavior (target) before training, making inference-time adaptability impossible. Given the large output space of object detection (including object existence prediction, bounding box estimation, and classification), the feasibility of flexible, inference-time model control remains unexplored. This paper introduces AnywhereDoor, a multi-target backdoor attack for object detection. Once implanted, AnywhereDoor allows adversaries to make objects disappear, fabricate new ones, or mislabel them, either across all object classes or specific ones, offering an unprecedented degree of control. This flexibility is enabled by three key innovations: (i) objective disentanglement to scale the number of supported targets; (ii) trigger mosaicking to ensure robustness even against region-based detectors; and (iii) strategic batching to address object-level data imbalances that hinder manipulation. Extensive experiments demonstrate that AnywhereDoor grants attackers a high degree of control, improving attack success rates by 26% compared to adaptations of existing methods for such flexible control.
Figure 1: Once implanted with AnywhereDoor, attackers can control the victim model with an unprecedented degree of freedom.
@@ 1 Introduction
Deep neural networks (DNNs) have revolutionized object detection [39, 37, 1], powering applications across autonomous vehicles [19], surveillance systems [24], medical imaging [25], and beyond [5, 27, 40]. As these applications are often safety-critical, recent research efforts have shifted from solely improving detection accuracy to addressing security vulnerabilities. Among the various threats, backdoor attacks are considered particularly serious in the industry [26]. In such attacks, a victim model appears to function normally until a secret pattern (e.g., a small white patch) is presented, causing the model to intentionally misbehave.
@@
However, existing studies on backdoor vulnerabilities in object detection are limited in scope. They assume a static, highly restrictive scenario in which the attacker predefines a single target and implants a corresponding trigger (e.g., using a small white patch to make any nearby object disappear) [29]. It remains unknown whether it is possible to design a backdoor that enables attackers to dynamically control the model based on context at inference time. While an intuitive solution is to extend existing methods by implanting multiple triggers, one for each possible target [2, 33, 10, 16], as we will show, this approach is insufficient in object detection due to its vast output space, which involves detecting a variable number of objects along with their bounding boxes and class labels [9, 8]. These limitations raise an intriguing question: Do backdoor attacks on object detection always have to be constrained to one or a small number of predefined targets?
In this paper, we answer this question by introducing AnywhereDoor, a multi-target backdoor attack that enables dynamic control over the victim model’s behavior. As illustrated in Figure 1, once implanted, AnywhereDoor allows an attacker to selectively make objects disappear, fabricate new ones, or mislabel them, either across all object classes or specific ones. The exact attack target can be chosen by the attacker based on the context at inference time to, e.g., maximize fatality. Our design is input-agnostic: the same trigger can be applied to any input (e.g., real-time video streams) to achieve consistent attacks.
Challenges. Achieving such control flexibility is challenging, especially in object detection. First, object detectors are multi-task learners, resulting in a vast number of possible attack targets. As shown in Figure 2, classic backdoor attacks like BadNet [21] (orange) may work fine when only one attack target (making “person” objects vanish) has to be supported. When more distinct triggers are injected to cover different targets (e.g., making “bus” objects vanish, misclassifying “person” objects as a “bottle,” etc.), its success rate drops drastically. Even for the state-of-the-art multi-target attack, Marksman [16] (blue), tested in image classification, it shows a better scalability in supporting more targets, but its effectiveness in object detection falls short. Second, many object detection algorithms adopt a region-based approach [49], extracting localized sub-regions of an image for object recognition. The trigger pattern can be corrupted (e.g., cropped) during this region proposal process [39] and fail to interfere with the prediction outcome. Third, object detection datasets often suffer from object-level imbalance (e.g., PASCAL VOC [17] contains 21×21$\times$21 × more instances of “person” than “bus”), which can lead to biased manipulability.
Figure 2: Existing backdoor methods like Marksman [16] and BadNet [21] are either not effective enough in object detection or are not scalable to support a large number of attack targets. In contrast, AnywhereDoor remains effective even when granting the attacker a high degree of control.
Contributions. To overcome the above challenges, we make four original contributions: (1) We present AnywhereDoor, the first backdoor attack emphasizing flexible control over object detection models; (2) We propose objective disentanglement, which scales the attack to support an unprecedented number of targets (see the red line in Figure 2) by factorizing and composing attack effects; (3) We introduce trigger mosaicking to promote the broad applicability across single-stage and two-stage detection paradigms, ensuring the trigger to remain effective within sub-regions of an image; (4) We propose strategic batching to redistributes learning opportunities in extremely imbalanced datasets common in object detection. Extensive experiments across various object detection algorithms and datasets confirm AnywhereDoor’s effectiveness, achieving over $26\%$ improvement in attack success rate while preserving the performance on clean samples.
@@ 2 Related Work
Backdoor Attacks on Image Classification. Pioneered by BadNet [21], backdoor attacks exploit the excessive learning capacity of DNNs [29] to associate a hidden trigger with a specific output by modifying a portion of the training data to (i) attach the trigger and (ii) alter the ground-truth label. Extensive efforts have been devoted to enhancing the stealthiness of backdoor attacks by designing invisible triggers in either the image domain [28, 15, 34] or the feature space [14, 6, 48], with some methods even avoiding label modification entirely [41, 43, 11]. Recently, multi-target attacks have gained attention [10, 23, 16], aiming to inject multiple carefully crafted triggers into the model to enable flexible control. However, directly applying these approaches to object detection yields low success rates due to its inherent limitations in backdooring object detectors and the large number of triggers that would need to be implanted.
Backdoor Attacks on Object Detection. The widespread use of object detection in safety-critical scenarios has motivated studies like BadDet [2] to investigate its resilience to backdoor attacks. These works predefine an attack target, such as making all or a certain class of objects vanished [7, 2, 4, 47, 33, 13], fabricated [7, 2, 4, 47], or mislabeled [4, 45, 13]. Current efforts focus on leveraging the unique properties of object detection to design more stealthy triggers, spanning both the physical [47, 13, 35, 33] and digital domains, including the use of natural object co-occurrence as a trigger [4]. Unlike prior work, which assumes a single target predefined before model training, this paper emphasizes the flexibility of backdoor attacks by enabling dynamic behavior alteration during inference.
@@ 3 Background
Threat Model. Consistent with prior multi-target backdoor attack settings [10, 23, 16], we consider the threat model, where the adversary has complete control of the training process of an object detector. Once the victim model is trained, it can be distributed through, e.g., model zoos for download by model users. During the inference phase, the adversary attempts to control the victim output by specifying the desired attack target, obtaining the trigger, and submitting the trigger-injected input to the victim model.
Figure 3: AnywhereDoor consists of three key modules: (a) objective disentanglement, which generates the trigger based on the attacker’s desired target; (b) trigger mosaicking, which overlays the trigger on a clean input by tiling to ensure its effectiveness in any subregions; and (c) strategic batching, which forms batches dynamically during training. Once trained, the victim object detector exhibits desired malicious behaviors when the trigger presents while outputting normal detection results when receiving a clean image as input.
Object Detection. Let $F_{\boldsymbol{\theta}}$ be a $K$-class object detection model parameterized by $\boldsymbol{\theta}$. Given an input image $\boldsymbol{x}$, the model returns a set of $N$ detected objects, denoted as $F_{\boldsymbol{\theta}}(\boldsymbol{x})=\hat{\boldsymbol{y}}=\{(\hat{\boldsymbol{b}}_{i},\hat{{c}}_{i})\}_{i=1}^{N}$, where $\hat{\boldsymbol{b}}_{i}$ represents the $i$-th detected bounding box, and $\hat{c}_{i}\in\{1,...,K\}$ denotes the corresponding predicted class label. Given a training dataset $\boldsymbol{\mathcal{D}}$ with pairs of input $\boldsymbol{x}$ and ground-truth objects $\boldsymbol{y}$, the model is trained by the following optimization problem:
$\boldsymbol{\theta}^{*}=\arg\min_{\boldsymbol{\theta}}\mathop{\mathbb{E}}_{(\boldsymbol{x},\boldsymbol{y})\in\boldsymbol{\mathcal{D}}}\mathcal{L}(\boldsymbol{y},F_{\boldsymbol{\theta}}(\boldsymbol{x}))$ (1)
where $\mathcal{L}$ is a loss function that quantifies the discrepancy between the detected and ground-truth objects. The specific formulation of it depends on the underlying object detection algorithm.
@@ 4 Methodology
Overview. AnywhereDoor automates the trigger design process by jointly optimizing the victim object detector $F_{\boldsymbol{\theta}}$ with a trigger generator $G_{\boldsymbol{\phi}}$, which takes the attack target as input and learns to generate the most effective trigger for manipulating the victim. Figure 3 gives an overview of AnywhereDoor’s training process, where it adopts a dirty-label data poisoning paradigm. Each training iteration consists of the following sequential operations:
  1. We strategically form a minibatch from the training dataset. To overcome the issues incurred by extreme data imbalances in object detection, we introduce a strategic batching to be presented in Section 4.3.
  2. We poison a portion of samples in the minibatch. For each of those training samples $(\boldsymbol{x},\boldsymbol{y})$, we (i) select a random attack target $\boldsymbol{e}$ from a pool of targets to be supported, (ii) send the attack target to the training-in-progress trigger generator $G_{\boldsymbol{\phi}}$ to generate the corresponding trigger $G_{\boldsymbol{\phi}}(\boldsymbol{e})$, (iii) attach the trigger to the input, and (iv) alter the ground-truth objects according to the selected attack target $\mathcal{P}(\boldsymbol{y};\boldsymbol{e})$. To scale to a large number of attack targets and promote the broad applicability to different object detection algorithms, we introduce an objective disentanglement and a trigger mosaicking design to be presented in Section 4.1 and Section 4.2, respectively.
  3. We use the standard object detection loss function (i.e., $\mathcal{L}$ in Equation 1) to optimize both the victim model $F_{\boldsymbol{\theta}}$ and the trigger generator $G_{\boldsymbol{\phi}}$ on the poisoned minibatch.
The joint optimization of $F_{\boldsymbol{\theta}}$ and $G_{\boldsymbol{\phi}}$ helps the trigger generator learn to create effective triggers for different attack targets, while the victim model learns to embrace being manipulated by those triggers.
Attack Scenarios. The motivation of AnywhereDoor is to create a backdoored object detector that can be manipulated with a wide range of attack targets. We generalize targets to five attack scenarios and provide their ground-truth alteration for a training sample $(\boldsymbol{x},\boldsymbol{y})$ as follows:
  • Untargeted removal deprives the victim model’s capability of recognizing the existence of any object (Figure 4(b)). For label poisoning, we eliminate all ground-truth objects (i.e., $\mathcal{P}(\boldsymbol{y};\boldsymbol{e})=\emptyset$)
  • Targeted removal only forces the victim to ignore objects of a given source class $c_{s}$ (Figure 4(c)). For label poisoning, we remove all ground-truth objects of class $c_{t}$ (i.e., $\mathcal{P}(\boldsymbol{y};\boldsymbol{e})=\{(\boldsymbol{b},c)\mid c\neq c_{s};(\boldsymbol{b},c)\in\boldsymbol{y}\}$). Note that by configuring different source classes, this attack scenario covers $K$ possible targets in a $K$-class object detection model, and we expect AnywhereDoor to support them all.
  • Untargeted misclassification requires the victim to still recognize object existence and their bounding boxes correctly, but all objects should be misclassified (Figure 4(d)). For label poisoning, we set the correct class label of each object to be the “next class,” according to the default indexing used in the original dataset (i.e., $\mathcal{P}(\boldsymbol{y};\boldsymbol{e})=\{(\boldsymbol{b},(c+1)\%K)\mid(\boldsymbol{b},c)\in\boldsymbol{y}\}$).
  • Targeted misclassification demands the victim to misclassify objects of a given source class $c_{s}$ to be a given destination class $c_{d}$ (Figure 4(e)). Other objects should be correctly detected. For label poisoning, we set the label of class $c_{s}$ objects to be $c_{d}$ (i.e., $\mathcal{P}(\boldsymbol{y};\boldsymbol{e})=\{(\boldsymbol{b},c_{d})\mid c=c_{s};(\boldsymbol{b},c)\in\boldsymbol{y}\}\cup\{(\boldsymbol{b},c)\mid c\neq c_{s};(\boldsymbol{b},c)\in\boldsymbol{y}\}$). Note that for a $K$-class object detection model, this attack scenario covers $(K-1)^{2}$ possible targets.
  • Untargeted generation fabricates fake objects to introduce false positives (Figure 4(f)). For label poisoning, all objects are duplicated with perturbations in their bounding box locations and sizes. In particular, for each duplicate of a ground-truth object $(\boldsymbol{b},c)\in\boldsymbol{y}$, we randomly choose a perturbation $\boldsymbol{\Delta}$ and apply it to the original bounding box (i.e., $\boldsymbol{b}+\boldsymbol{\Delta})$ while ensuring the perturbed object remains inside the image.
In the remainder of this section, we delve into the design of the key modules in AnywhereDoor to make the simultaneous support of all the above attack scenarios possible.
(a) Ground Truth Label
(b) Poisoned Label:
@@
(c) Poisoned Label:
@@
(d) Poisoned Label:
@@
(e) Poisoned Label:
@@
(f) Poisoned:
@@
Figure 4: Five attack scenarios (b-f) supported by AnywhereDoor with their poisoned labels on a training sample as an example.